# SpeechAct: Towards Generating Whole-body Motion from Speech

### v3

@@ 3.1.3 Translation Model
With the learned motion codebook, as shown in Fig. 2 (b), we aim to translate the audio signal into the motion representation from the motion codebook, and then output the motion sequence using the learned motion decoder in the second stage. Speech content, especially phonetic information, is closely related to face motion, notably lip motion, while its link to body/hand motion is weaker. Compared with phonetic information in the audio signal, speech rhythm and beat are more related to body and hand motion. Therefore, given the input audio signal $A=\{a_{t}\}_{t=0}^{T}$, we first transform it to 64-dimensional Mel-Frequency Cepstral Coefficients (MFCC) features $F_{a}$ containing rhythm and beat information. To regress the motion representation with speaker information, we design a translation model $\mathcal{T}$ to translate the input audio signal into the motion codes. Specifically, we concatenate the speaker embedding, i.e., a one-shot embedding, with $F_{a}$, and employ several retention blocks [53] to extract the regressed features $F_{r}\in\mathcal{R}^{\frac{T}{w_{m}}\times N}$, where $w_{m}$ is the temporal window size and $N$ is the length of the motion codebook. Subsequently, we adopt a SoftMax layer to obtain the regressed logits, and select the highest score at each temporal window as the index of the quantized feature. Thus, we can obtain the generated features $F_{g}$ by selecting features according to the selected indexes from the learned motion codebook. With the learned motion decoder, we can obtain reasonable results in harmony with the input audio. To regress reasonable results, we adopt cross-entropy loss $\mathcal{L}_{ce}$ to optimize the translation model.
@@
Contrastive Motion Learning. Because the input audio and the output motion are one-to-one mappings, the diversity of the generated results is still limited. An intuitive idea is that, given a speech recording, the motion for a specific speaker can be different from other speakers, and can be different from the motions driven by other speech recordings. To achieve this, we introduce a novel contrastive motion learning method, which can distinguish the current generated motion from other motions driven by different audios or generated by other speakers. As shown in Fig. 3, different from previous works [38, 8] using contrastive loss to align multi-modal information, e.g., text and motion, we take the early generated quantized features from other speakers and other speech recordings as the negative samples, and the ground-truth quantized features as the positive sample to boost the diversity of generated results. Specifically, we adopt global average pooling along time dimension to obtain motion codes as the negative or positive samples. Evidently, the positive sample is high-quality and in harmony with the input audio signal, while the negative samples are low-quality and have no connection with the input audio signal. Let $p$ denote the current generated features $F_{g}$, $p^{+}$ that of groud-truth quantized feature from $V$, and $\{p_{t}^{-}\}_{t=1}^{L}$ the generated features from $L$ negative samples. To boost the diversity of the generated results, our goal is to ensure that the generated features are different from the negative samples, i.e., minimizing mutual information. Therefore, the contrastive loss can be formulated as:
@@ 3.2 Face Generator
Because the face is closely related to the input signal, especially phonetic information, as shown in Fig. 2 (c), we adopt an encoder-decoder architecture to regress the facial movement. Specifically, we first encode the audio signal $A=\{a_{i}\}_{i=0}^{T}$ using a pre-trained WavLM model, where the feature extraction is adapted to generate a representation suitable for downstream tasks. This representation is concatenated with speaker embedding to obtain speaker information, and then refined by an adaptive module to extract and aggregate the audio features adaptively. The adaptive module consists of a residual block, followed by an attention block, and then two more residual blocks. This hierarchical arrangement of convolutional and attention mechanisms aims to robustly capture and understand the speaker-specific characteristics and the speech content to generate more reliable results. Subsequently, a series of residual blocks with attention mechanisms are used to decode the audio representation, transforming it into a deterministic face motion sequence. The detailed architecture can be found in the supplemental document. The face generator is trained with face reconstruction loss: L1 distance for expressions and L1 distance for the first order of the face motion representations.
3.3 Training Details
We conduct our experiments on a desktop with a GeForce RTX 3090 GPU. For the face generator, we adopt the parametric representation for deterministic results, and the parameters of WavLM are frozen. For the body generator, we adopt our proposed hybrid points representation. We first train the VQ-VAE to learn the motion codebook, and the generator from our representations to obtain parameters of SMPL-X. Then, we train the translation model and freeze the parameters of the motion decoder. For all training models, we adopt the Adam [54] optimizer with coefficients 0.9 and 0.999 for computing running averages of gradients and their squares. The learning rates for different stages are all set to 1e-4.
@@ 4.1 Experimental Setting
Baselines. To validate the performance of our model, We first compare with TalkShow [5], which is the most related work that generates whole-body motion from speech. Besides, we also compare with three state-of-the-art methods: Audio2Gesture [32], DiffStyle [39] and EMAGE [42]. Because Audio2Gesture and TalkShow are trained using other part body dataset, we re-train them on BEAT2 dataset using the public codes. Besides, we implement two baselines for generating body and hand motions to validate the effectiveness of our model. All the experiments are conducted on the same training and testing set. It should be noted that DiffStyle and EMAGE are tested on 1 speaker. Therefore, we adopt the same speaker to evaluate their performance additionally for a fair comparison with them. Additionally, we implement two baselines to validate the effectiveness of our model.
@@
  • Motion VAE. In this model, we first train a VAE using motion sequences to learn a latent space. Then, given a speech, we encode the audio signal and concatenate it with a sampling latent code from latent space to output the generated motion using a decoder.
Metric. For the body and hands, we aim to generate reasonable and diverse results in harmony with the input audio. Therefore, we adopt several evaluation metrics to validate the diversity and realism of generated results.
@@
  • Div-in: Following [57], we evaluate the diversity of motions corresponding to an individual speech through variations in body poses across the temporal sequence, which is denoted as Div-in.
  • Div-out: Given a specific speech, diverse results for the same speaker are expected. To evaluate this, for all audio clips in our test set, we calculate the average L2 distance between the two motion clips generated from an audio clip, denoted as Div-out.
TABLE I: Quantitative comparison of body motion on BEAT2 dataset with several baselines. Ours means we evaluate our method on the single speaker for a fair comparison with DiffStyle and EMAGE.
Audio Enc-Dec 10.36 8.907 0.5504 0.8632 0
Motion VAE 10.73 8.722 0.5363 0.8733 0.0571
Audio2Gesture [32] 20.43 4.826 0.5572 0.3893 1.312
TalkShow [5] 14.37 4.535 0.5479 0.8966 1.744
Ours 3.839 4.359 0.5603 0.9700 2.264
DiffStyle [39] 44.91 6.576 0.5811 1.539 0.3340
EMAGE [42] 6.620 10.12 0.5487 2.817 0
Ours 5.979 5.505 0.5704 3.335 7.840
@@
Figure 5: Visualization of the diversity of different speakers with different speeches and the diversity of a single speaker with different speeches compared with three state-of-the-art methods. The inter-class distance of each method is shown in the top right corner of each figure. Different colors in (a) indicate the visualization of the generated results conditioned on different speakers, and each speaker in the same method has a specific speech. Different colors in (b) show the visualization of different speeches talked by the same speaker.
@@ 4.2 Comparison Results
Quantative Evaluation. For body and hand motion generation, Table I presents the quantitative comparison on BEAT2 dataset with several baselines. It can be seen that our method outperforms other methods in most metrics. Specifically, our method significantly improves the performance on FID (both FID-k and FID-g), which indicates that our model generates more realistic body and hand motion sequences. This is because our hybrid point representation can help our model learn a better motion codebook in quantized motion modeling, which is important to generate reasonable results. Also, our method achieves the best performance on BeatAlign score. This denotes that the motion sequences generated by our method are more rhythm-consistent with the input audio. Moreover, the diversity of our results outperforms other baselines, which indicates our contrastive motion learning method can generate more distinctive motion representations.
@@
Qualitative Evaluation. Fig. 4 shows the qualitative comparison compared with TalkShow [5] and EMAGE [42]. In the left subfigure, each sample consists of five frames extracted at intervals of 2/15 seconds from a generated motion clip. It can be seen that the foot movements generated by TalkShow and EMAGE change rapidly in a short time, leading to foot skating. In contrast, due to our hybrid point representation, our model can generate continuous and smooth motion movements. Besides, our method generates realistic movements corresponding to the audio input, e.g., hands down when emphasizing “after all”. The right subfigure in Fig. 4 presents two generated results given two audio inputs. It can be seen that our method can generate more diverse body and hand motions for a single sentence, which benefits from our contrastive motion learning method. The dynamic results can be found in the supplementary video.
To boost the diversity of the generated results, we design a contrastive motion learning method according to an intuitive idea that the generated motion of specific audio should be different from the generated motions of other speakers and other audios. To verify the performance of our model, we visualize generated results using tSNE method to present the diversity of different speakers with different speeches and the diversity of a single speaker with different speeches. Fig. 5 illustrates the visualization of the generated results. We also give the inter-class distance ($s$) to present the diversity between the different clusters. The inter-class distance refers to the measure of separation between different groups. Fig. 5 (a) is obtained from 8 speakers with different speeches, which indicates that the generated motions should be different. It can be seen that our model has better performance, which can generate more distinctive motions different from other speakers. On the contrary, the other methods struggle to generate distinctive motions, leading to similar motions even among different speakers with different sentences. This is also the reason that the inter-class distance of TalkShow (23.5) is closer to our method (25.1). Fig. 5 (b) illustrates the diversity of a single speaker with different speeches. We can see that with different speeches, our method achieves better cluster performance and gets the best inter-class distance, which indicates that our method can generate more diverse results for a single speaker.
Figure 6: Visualization of reconstructed motion errors. The errors between the reconstructed motions and the ground-truths are color-coded on the reconstructed models for visual inspection
@@
Representation Keypoints Vertices
@@
3D (TalkShow) 0.4409 14.42 26.71 10.87
6D 0.4514 15.76 26.98 11.86
Points 0.3095 10.85 21.68 7.896
Keypoints 0.2836 10.68 19.39 8.595
Ours 0.2744 7.936 16.68 6.050
@@ 4.3 Ablation Study
Effectiveness of hybrid point representation. To validate the effectiveness of our hybrid point representation, instead of conducting experiments on the motion generation task, we compare different representations on the motion reconstruction task. Specifically, we replace our representation with different representations for the first stage of our body generator, i.e., quantized motion modeling (Sec. 3.1.2). We employ the 3D representation, 6D representation, surface points of SMPL-X body mesh, and keypoints of SMPL-X body mesh. The model settings are the same as our model. It should be noted that the model with 3D representation has the same representation as TalkShow.
We adopt landmark velocity difference (LVD) and landmark distance (LD) to evaluate the quality of reconstructed motion. To evaluate different representations, we extract keypoints and vertices of reconstructed SMPL-X body meshes to compute LVD and LD metrics. Table III presents the quantitative results on BEAT2 dataset for different motion representations. It can be seen that compared with the point representation, the parametric representation can generate inaccurate results (high LD score) with jitters (high LVD score). Comparing points representation and keypoints representation, our hybrid representation can achieve more accurate and smoother results, which indicates that our representation is suitable as a motion representation. Besides, this is also the reason that our model achieves better results in Table I compared with other methods.
Fig. 6 shows the qualitative results. It can be seen that the models based on 3D and 6D representations can produce errors for local body parts, such as hands and legs. The keypoint-based model has better performance, however, the arms are difficult to recover due to a lack of global constraints of the pose rotation. The point-based model can recover more accurate arms, but it struggles to reconstruct the head and legs accurately due to redundant information for some body parts. Compared with these traditional representations, our hybrid point representation can recover the body accurately, which indicates its strong capacity for motion reconstruction. With this powerful representation, we can learn a good motion codebook to further help generate natural and reasonable motions from speech.
@@
Figure 7: Visualization of the diversity of different speakers with different speeches (a) and the diversity of a single speaker with different speeches (b). The inter-class distance is shown at the top left corner of each figure.
@@
Fig. 7 shows the visualization of motions generated by our model. Compared with w/o CM and our full model, it can be seen that the contrastive learning method can help the model to generate more distinctive motions for different speakers and different speeches. The increase in the inter-class distances of our full model also verifies this.

### v5

@@ 3.1.3 Translation Model
With the learned motion codebook, as shown in Fig. 2 (b), we aim to translate the audio signal into the motion representation from the motion codebook, and then output the motion sequence using the learned motion decoder in the second stage. Speech content, especially phonetic information, is closely related to face motion, notably lip motion, while its link to body/hand motion is weaker. Compared with phonetic information in the audio signal, speech rhythm and beat are more related to body and hand motion. Therefore, given the input audio signal $A=\{a_{t}\}_{t=0}^{T}$, we first transform it to 64-dimensional Mel-Frequency Cepstral Coefficients (MFCC) features $F_{a}$ containing rhythm and beat information. To regress the motion representation with speaker information, we design a translation model $\mathcal{T}$ to translate the input audio signal into the motion codes. Specifically, we concatenate the speaker embedding, i.e., a one-shot embedding, with $F_{a}$, and employ several retention blocks [57] to extract the regressed features $F_{r}\in\mathcal{R}^{\frac{T}{w_{m}}\times N}$, where $w_{m}$ is the temporal window size and $N$ is the length of the motion codebook. Subsequently, we adopt a SoftMax layer to obtain the regressed logits, and select the highest score at each temporal window as the index of the quantized feature. Thus, we can obtain the generated features $F_{g}$ by selecting features according to the selected indexes from the learned motion codebook. With the learned motion decoder, we can obtain natural results in harmony with the input audio. To regress natural results, we adopt cross-entropy loss $\mathcal{L}_{ce}$ to optimize the translation model. As an auto-regressive model, the translation model can generate motions of arbitrary length by adjusting the input sequence of previously predicted tokens. Besides, during inference, we randomly initialize the first token using the index of the codebook to generate diverse results.
@@
Contrastive Motion Learning. Because the input audio and the output motion are one-to-one mappings in training pairs, the diversity of the generated results is still limited. An intuitive idea is that, given a speech recording, the motion for a specific speaker can be different from other speakers, and can be different from the motions driven by other speech recordings. To achieve this, we introduce a novel contrastive motion learning method, which can distinguish the current generated motion from other motions driven by different audios or generated by other speakers. As shown in Fig. 4, different from previous works [42, 9] using contrastive loss to align multi-modal information, e.g., text and motion, we take the early generated quantized features from other speakers and other speech recordings as the negative samples, and the ground-truth quantized features as the positive sample to boost the diversity of generated results. Specifically, we adopt global average pooling along time dimension to obtain motion codes as the negative or positive samples. Evidently, the positive sample is high-quality and in harmony with the input audio signal, while the negative samples are low-quality and have no connection with the input audio signal. Let $p$ denote the current generated features $F_{g}$, $p^{+}$ that of groud-truth quantized feature from $V$, and $\{p_{t}^{-}\}_{t=1}^{L}$ the generated features from $L$ negative samples. To boost the diversity of the generated results, our goal is to ensure that the generated features are different from the negative samples, i.e., minimizing mutual information. Therefore, the contrastive loss can be formulated as:
@@ 3.2 Face Generator
In this work, our goal is to generate whole-body motion from speech audio, and the most important part is body generation due to its diverse character. To support the overall whole-body motion synthesis, we include a face generator to generate facial motion from speech audio. Because the face is closely related to the input signal, especially phonetic information, as shown in Fig. 2 (c), we adopt an encoder-decoder architecture to regress the facial movement. Specifically, we first encode the audio signal $A=\{a_{i}\}_{i=0}^{T}$ using a pre-trained WavLM model [58], where the feature extraction is adapted to generate a representation suitable for downstream tasks. This representation is concatenated with speaker embedding to obtain speaker information, and then refined by an adaptive module to extract and aggregate the audio features adaptively. The adaptive module consists of a residual block, followed by an attention block, and then two more residual blocks. This hierarchical arrangement of convolutional and attention mechanisms aims to robustly capture and understand the speaker-specific characteristics and the speech content to generate more reliable results. Subsequently, a series of residual blocks with attention mechanisms are used to decode the audio representation, transforming it into a deterministic face motion sequence. The detailed architecture can be found in the supplemental document. It should be noted that the face generator serves as a complementary aspect and can be replaced by state-of-the-art facial motion synthesis models [59, 60, 61, 62]. The face generator is trained with face reconstruction loss: L1 distance for expressions and L1 distance for the first order of the face motion representations.
3.3 Implemention Details
We conduct our experiments on a desktop with a GeForce RTX 3090 GPU. For the face generator, we adopt the parametric representation for deterministic results, and the parameters of WavLM are frozen. For the body generator, we adopt our proposed hybrid points representation. We first train the VQ-VAE to learn the motion codebook, and the generator from our representations to obtain parameters of SMPL-X. The weights ${\alpha_{1},\alpha_{2},\beta}$ of different losses are set to $0.5,0.5,0.25$. We set the channel dimensions of quantized features $Z_{q}^{lh},Z_{q}^{rh},Z_{q}^{b}$ as $128,128,512$. The length of the temporal window size is set to $\frac{1}{30}$ s for each input audio. Then, we train the translation model and freeze the parameters of the motion decoder. For all training models, we adopt the Adam [63] optimizer with coefficients 0.9 and 0.999 for computing running averages of gradients and their squares. The learning rates for different stages are all set to 1e-4.
@@ 4.1 Experimental Setting
Baselines. To validate the performance of our model, We first compare with TalkShow [1], which is the most related work that generates whole-body motion from speech. Besides, we also compare with three state-of-the-art methods: Audio2Gesture [36], DiffStyle [2] and EMAGE [45]. Because Audio2Gesture and TalkShow are trained using other part body dataset, we re-train them on BEAT2 dataset using the public codes. Besides, we implement two baselines for generating body and hand motions to validate the effectiveness of our model. All the experiments are conducted on the same training and testing set. It should be noted that DiffStyle and EMAGE are tested on 1 speaker. Therefore, we adopt the same speaker to evaluate their performance additionally for a fair comparison with them.
@@
  • Motion VAE. In this model, we first train a VAE using motion sequences to learn a latent space. Then, given a speech audio, we encode the audio signal and concatenate it with a sampling latent code from latent space to output the generated motion using a decoder.
Metrics. For the body and hands, we aim to generate natural and diverse results in harmony with the input audio. Therefore, we adopt several evaluation metrics to validate the diversity and realism of generated results.
@@
  • Div-in: Following [66], we evaluate the diversity of motions corresponding to an individual speech audio through variations in body poses across the temporal sequence, which is denoted as Div-in.
  • Div-out: Given a specific speech audio, diverse results for the same speaker are expected. To evaluate this, for all audio clips in our test set, we calculate the average L2 distance between the two motion clips generated from an audio clip, denoted as Div-out.
  • FSR: With our hybrid representation, our model can generate smooth results alleviating foot skating. To validate this character, following [67, 44], we identify frames where the foot slides beyond a certain distance while in contact with the ground (i.e., foot height <5 cm) as foot-skating frames. We report the Foot Skating Ratio (FSR), which measures the proportion of such frames.
TABLE I: Quantitative comparison of body motion on BEAT2 dataset with several baselines. Ours means we evaluate our method on the single speaker for a fair comparison with DiffStyle and EMAGE. Note that we report FSR ×$\times 10^{-2}$.
Model FID-k $\downarrow$ FID-g $\downarrow$ BeatAlign $\uparrow$ Div-in $\uparrow$ Div-out $\uparrow$ FSR $\downarrow$
Audio Enc-Dec 10.36 8.907 0.5504 0.8632 0 2.99
Motion VAE 10.73 8.722 0.5363 0.8733 0.0571 6.78
Audio2Gesture 20.43 4.826 0.5572 0.3893 1.312 1.26
TalkShow 14.37 4.535 0.5479 0.8966 1.744 0.61
Ours 3.839 4.359 0.5603 0.9700 2.264 0.38
DiffStyle 44.91 6.576 0.5811 1.539 0.3340 1.69
EMAGE 6.620 10.12 0.5487 2.817 0 2.28
Ours 5.979 5.505 0.5704 3.335 7.840 1.19
@@
Figure 7: Visualization of the diversity of different speakers with different speech audios and the diversity of a single speaker with different speech audios compared with three state-of-the-art methods. The inter-class distance of each method is shown in the top right corner of each figure. Different colors in (a) indicate the visualization of the generated results conditioned on different speakers, and each speaker in the same method has a specific speech audio. Different colors in (b) show the visualization of different speech audios talked by the same speaker.
@@ 4.2 Comparison Results
Quantative Evaluation. For body and hand motion generation, Table I presents the quantitative comparison on BEAT2 dataset with several baselines. It can be seen that our method outperforms other methods in most metrics. Specifically, our method significantly improves the performance on FID (both FID-k and FID-g), which indicates that our model generates more realistic body and hand motion sequences. The metric FID-k focuses on the speed and acceleration of movement, reflecting the physical realism of the generated motions. Results generated by previous often contain foot-skating frames and disharmony contacts between ground and feet, which causes unnatural speed and acceleration of the movements. While our method can generate more natural and smooth motion sequences due to our hybrid representation, which leads to the high improvements on FID-K. Besides, the comparison on FSR demonstrates that our model can generate more smooth motions and can alleviate foot skating in generated motions, which suggests the effectiveness of our hybrid representation. Also, our method achieves the best performance on BeatAlign score. This denotes that the motion sequences generated by our method are more rhythm-consistent with the input audio. Moreover, the diversity of our results outperforms other baselines, which indicates our contrastive motion learning method can generate more distinctive motion representations. The diffusion-based models, i.e., DiffStyle, requires a large number of diffusion steps, leading to 7 times longer than our method to generate a 4-second motion sequence (5.80s vs. 0.78s). The detailed comparison on inference time can be found in the supplemental document.
@@
Qualitative Evaluation. Fig. 5 shows the qualitative comparison compared with TalkShow [1] and EMAGE [45]. In the left subfigure, each sample consists of five frames extracted at intervals of 2/15 seconds from a generated motion clip. It can be seen that the foot movements generated by TalkShow and EMAGE change rapidly in a short time, leading to foot skating. In contrast, due to our hybrid point representation, our model can generate continuous and smooth motion movements. Besides, our method generates realistic movements corresponding to the audio input, e.g., hands down when emphasizing “after all”. The right subfigure in Fig. 5 presents two generated results given two audio inputs. It can be seen that our method can generate more diverse body and hand motions for a single sentence, which benefits from our contrastive motion learning method. Fig. 6 presents the comparison of generated facial movements. The results show that our method can generate more accurate motions for both open-mouth and closed-mouth sounds. More dynamic results and analyses of different methods can be found in the supplementary material. We also provide some long generation samples (over 3 minutes) in the supplemental video.
To boost the diversity of the generated results, we design a contrastive motion learning method according to an intuitive idea that the generated motion of specific audio should be different from the generated motions of other speakers and other audios. To verify the performance of our model, we visualize generated results using tSNE method to present the diversity of different speakers with different speech audios and the diversity of a single speaker with different speech audios. Fig. 7 illustrates the visualization of the generated results. We also give the inter-class distance ($s$) to present the diversity between the different clusters. The inter-class distance refers to the measure of separation between different groups. Fig. 7 (a) is obtained from 8 speakers with different speech audios, which indicates that the generated motions should be different. It can be seen that our model has better performance, which can generate more distinctive motions different from other speakers. On the contrary, the other methods struggle to generate distinctive motions, leading to similar motions even among different speakers with different sentences. This is also the reason that the inter-class distance of TalkShow (23.5) is closer to our method (25.1). Fig. 7 (b) illustrates the diversity of a single speaker with different speech audios. We can see that with different speech audios, our method achieves better cluster performance and gets the best inter-class distance, which indicates that our method can generate more diverse results for a single speaker.
Figure 8: Visualization of reconstructed motion errors. The errors between the reconstructed motions and the ground-truths are color-coded on the reconstructed models for visual inspection. Each row shows reconstructed error for each case.
@@
Representation Keypoints Vertices Param.
@@
3D (TalkShow) 0.4409 14.42 26.71 10.87 82.4 M
6D 0.4514 15.76 26.98 11.86 82.7 M
Surface Points 0.3095 10.85 21.68 7.896 88.5 M
Keypoints 0.2836 10.68 19.39 8.595 84.0 M
Ours 0.2744 7.936 16.68 6.050 89.4 M
@@ 4.3 Ablation Study
Effectiveness of hybrid point representation. To validate the effectiveness of our hybrid point representation, instead of conducting experiments on the motion generation task, we compare different representations on the motion reconstruction task. Specifically, we replace our representation with different representations for the first stage of our body generator, i.e., quantized motion modeling (Sec. 3.1.2). We employ the 3D representation, 6D representation, surface points of SMPL-X body mesh, and keypoints of SMPL-X body mesh. Note that aside from the dimensions of the input and output layers, all other layer configurations, such as the number of layers and hidden dimensions, are identical. It should be noted that the model with 3D representation has the same representation as TalkShow.
We adopt landmark velocity difference (LVD) and landmark distance (LD) to evaluate the quality of reconstructed motion. To evaluate different representations, we extract keypoints and vertices of reconstructed SMPL-X body meshes to compute LVD and LD metrics. In addition, we report the number of parameters (Param.) for different representations. Table III presents the quantitative results on BEAT2 dataset for different motion representations. It can be seen that compared with the surface point representation, the parametric representation can generate inaccurate results (high LD score) with jitters (high LVD score). It is interesting that the performance of 6D representation is lower than 3D representation. The main reason is that speech motion, involves a lot of local rotations, which typically involve small angle changes. Therefore, the 3D axis-angle representation might more easily capture these local characteristics, leading to better performance in the reconstruction task. Comparing surface points representation and keypoints representation, our hybrid representation can achieve more accurate and smoother results, which indicates that our representation is suitable as a motion representation. This is also the reason that our model achieves better results in Table I compared with other methods.
Fig. 8 shows the qualitative results. It can be seen that the models based on 3D and 6D representations can produce errors for local body parts, such as hands and legs. The keypoint-based model has better performance, however, the arms are difficult to recover due to a lack of global constraints of the pose rotation. The point-based model can recover more accurate arms, but it struggles to reconstruct the head and legs accurately due to redundant information for some body parts. Compared with these traditional representations, our hybrid point representation can recover the body accurately, which indicates its strong capacity for motion reconstruction. With this powerful representation, we can learn a good motion codebook to further help generate natural motions from speech audios.
@@
Fig. 9 shows the visualization of motions generated by our model. Compared with w/o CM and our full model, it can be seen that the contrastive learning method can help the model to generate more distinctive motions for different speakers and different speech audios. The increase in the inter-class distances of our full model also verifies this.
Effectiveness of separate codebooks. We use three codebooks, i.e., the body codebook, the left-hand codebook, and the right-hand codebook, to model more detailed motions for the body and hands, respectively. To validate the effectiveness of this design, we conduct experiments using different numbers of codebooks. Specifically, we train motion reconstruction models using one codebook, two codebooks (hand and body codebooks), and three codebooks (our current design). Table V illustrates the quantitative results. As shown, increasing the number of codebooks improves the performance of motion reconstruction. Although the landmark velocity distance (LVD) for joints is better with two codebooks, the other metrics are lower compared to the model with three codebooks. Therefore, we adopt three separate codebooks in our model.
Figure 9: Visualization of the diversity of different speakers with different speech audios (a) and the diversity of a single speaker with different speech audios (b). The inter-class distance is shown at the top left corner of each figure.
TABLE V: Ablation studies on BEAT2 dataset for different numbers of codebooks. Note that we report LD×$\times 10^{2}$ for vertices.
Number Keypoints Vertices
LVD $\downarrow$ LD $\downarrow$ LVD $\downarrow$ LD $\downarrow$
1 0.2776 8.098 16.78 6.226
2 0.2683 8.002 16.71 6.179
3 0.2744 7.936 16.68 6.050